OpenStudy (anonymous):

Can someone Help Me? D: Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (johnweldon1993):

Well f(1) is 2 f(2) = 4 f(n) = f(1) + f(2) + f(n-1) So if we have f(3) we have f(1) + f(2) + f(3-1) or f(1) + f(2) + f(2) = 2 + 4 + 4 = 10 for f(4) it would be f(1) + f(2) + f(4-1) or f(1) + f(2) + f(3) = 2 + 4 + 10 = 16 Finally f(5) we have f(1) + f(2) + f(5-1) or f(1) + f(2) + f(4) = 2 + 4 + 16 = 22 f(5) = 22